Data from an experiment show that the standard deviation is 3.6 and the mean is 23.4.
Find the probability that the reading from the experiment is more than 32.
1
Expert's answer
2020-10-04T17:29:34-0400

Given that,


Standard deviation "\sigma = 3.6"

Population mean "\mu" = 23.4

experiment reading x = 32


We have to find the probability of greater than Z"\alpha" if x>32


Formula for Z"\alpha" = "\frac {x-\mu}{\sigma} = \frac {32-23.4}{3.6} = 2.3888"


So by using the Z table we get the probability of greater than Z"\alpha" if x>32 is 1 - 0.9916 = 0.0084.